Can someone please help me with this question.
Calculate the number of male and female students required to ensure a sample of 10 students contains the same proportions as the full population, which contains 22 male and 28 female students.
Thanks alot

Answer by stanbon(48568)

Calculate the number of male and female students required to ensure a sample of 10 students contains the same proportions as the full population, which contains 22 male and 28 female students.
male/female = 22/28=11/14
A sample of size 10 is not large enough to have whole human beings
in the desired proportion.
A sample of size 100 would do it.
The numbers would be 44 male and 56 female.
You get this by considering the whole number multiples of 11 and 14
4*11 = 44 and 4*14= 56 and 44+56 = 100
Cheers,
Stan H.